# KCL-PH-TH/2025-19

### v1

@@ Abstract
We investigate two classes of dark matter (DM) candidates, sub-GeV particles and primordial black holes (PBHs), that can inject low-energy electrons and positrons into the Milky Way and leave observable signatures in the X-ray sky. In the case of sub-GeV DM, annihilation or decay into $e^{+}e^{-}$ contributes to the diffuse sea of cosmic-ray (CR) leptons, which can generate bremsstrahlung and inverse Compton (IC) emission on Galactic photon fields, producing a broad spectrum from X-rays to $\gamma$-rays detectable by instruments such as eROSITA and XMM-Newton. For PBHs with masses below $\sim 10^{17}$ g, Hawking evaporation similarly yields low-energy $e^{\pm}$, leading to comparable diffuse emission. Using the first data release from eROSITA and incorporating up-to-date CR propagation and diffusion parameters, we derive new constraints on both scenarios. For sub-GeV DM, we exclude thermally averaged annihilation cross sections in the range $\sim 10^{-27}$$10^{-25}\ \mathrm{cm^{3}/s}$ and decay lifetimes of $\sim 10^{24}$$10^{25}$ s for masses between 1 MeV and 1 GeV, with eROSITA outperforming previous X-ray constraints below $\sim 30$ MeV. For asteroid-mass PBHs, we set new bounds on the DM fraction based on their Hawking-induced emission. Finally, we revisit earlier constraints from XMM-Newton, finding that they were approximately four orders of magnitude too stringent due to the use of the instrument’s geometric solid angle rather than its exposure-weighted solid angle. Upon using the exposure-weighted solid angle, we show that the revised XMM-Newton limits are slightly weaker than those from eROSITA.
@@ 1 Introduction
In this work, we address this gap by placing new constraints on light DM and sub-asteroid mass PBHs using eROSITA’s full-sky X-ray maps. We model continuum emission from DM-induced low-energy $e^{\pm}$ and Hawking evaporation from PBHs, deriving upper limits on the annihilation cross section, decay lifetime, and PBH DM fraction across a wide mass range. We then compare these constraints with existing bounds from other probes, demonstrating the complementarity and reach afforded by eROSITA. Additionally, we revisit constraints derived from XMM-Newton data [27], showing that applying the geometric solid angle instead of the exposure-weighted solid angle led to flux differences of approximately four orders of magnitude. After addressing this, the XMM-Newton limits are found to be weaker than those from eROSITA, realigning the landscape of X-ray DM constraints.
@@ 3.1 Sub-GeV dark matter annihilation and decay
We restrict our analysis to the channel of direct $e^{+}e^{-}$ production, from a minimum kinematically allowable mass of $m_{e}$ ($2m_{e}$) for annihilation (decay) respectively. In both cases, we take the maximum mass of DM to be $1000$ MeV. We use a 2D axisymmetric Galactic model with a NFW DM profile [36] representing our base case. Other decay channels including the production of pions and muons can also be considered, but the constraints become much weaker than in previous analyses for the high DM masses needed. The propagation parameters used in this work are obtained from a fit to AMS-02 data [37, 38], and are listed in Table I of Ref. [32], where more of the technical details are given. The interplay between diffusion and energy losses is key at sub-GeV energies. Energy losses in the DRAGON2 code include ionisation, Coulomb, bremsstrahlung, IC, synchrotron and in-flight positron annihilation processes. The spatial and energy resolution employed is $\sim 150$ pc and $5$%.
@@ 4 Constraints from eROSITA data
Finally, in the bottom panel of Fig. 3, we also report constraints on PBHs in the mass range $10^{14}\ \mathrm{g}\lesssim m_{\mathrm{PBH}}\lesssim 10^{17}\ \mathrm{g}$ from Hawking evaporation of charged particles (orange line). As commented above, these constraints are derived assuming monochromatic and non-rotating PBHs. The orange band again represent uncertainties arising from propagation. Here, we also compare our result from eROSITA with other leading constraints, including the 511 keV line, XMM-Newton and Voyager 1 (as green lines). In comparison to these probes, eROSITA diffuse background observations do not provide competitive constraints.
@@ 5 Revising XMM-Newton constraints
As noted in the previous section, the DM constraints derived from eROSITA data are significantly weaker than those reported in earlier analyses of XMM-Newton observations [14, 32]. At first glance, this appears counterintuitive. While eROSITA probes lower energy bands than XMM-Newton, where the latter suffers from increased instrumental noise, one would still expect comparable fluxes in the same regions of the sky when interpreted correctly.
We believe this discrepancy arises from a misinterpretation in previous analyses regarding the solid angle associated with the data presented in Ref. [27]. In particular, earlier works assumed the reported fluxes were distributed over the full geometric solid angle of each ring. However, the data in Ref. [27] are normalised to the exposure-weighted average solid angle $\Omega_{\mathrm{exp}}$, which accounts for the non-uniform exposure across the ring due to the telescope’s observational footprint [^2] A Jupyter notebook tutorial on how to interpret the data is now provided in the original Github repository.. This is defined
@@
Figure 5: Comparison of ring 3 flux for eROSITA (orange points) and XMM-Newton MOS with what the use of exposure-weighted solid angle (black points) consistent with observational coverage of the dataset and with direct use of the geometric solid angle (grey points). In order to enable a viable comparison, all fluxes shown here use a 2 mask of the Galactic plane, as provided for the XMM-Newton data.
The discrepancy in solid angle varies between rings, so a full reanalysis is required to derive revised XMM-Newton constraints. Using the same analysis framework as for eROSITA, we compute updated XMM-Newton limits on DM annihilation and decay into $e^{\pm}$, as well as PBH evaporation, shown in Fig. 6. These incorporate propagation uncertainties consistent with those used in the eROSITA analysis. We do not repeat the uncertainty propagation for the DM density distribution or the local DM density as this can be inferred easily from Fig. 4.
@@ 6 Discussion and conclusions
Finally, we reanalysed previous constraints from XMM-Newton data and identified a discrepancy arising from the treatment of the instrument’s exposure-weighted solid angle. Specifically, an overestimate of the effective sky coverage led to flux values, and therefore constraints, being overstated by approximately four orders of magnitude. After correcting this effect and rederiving the XMM-Newton limits using consistent assumptions, we find that they are weaker than those derived from eROSITA.

### v2

@@ Abstract
We investigate two classes of dark matter (DM) candidates, sub-GeV particles and primordial black holes (PBHs), that can inject low-energy electrons and positrons into the Milky Way and leave observable signatures in the X-ray sky. In the case of sub-GeV DM, annihilation or decay into $e^{+}e^{-}$ contributes to the diffuse sea of cosmic-ray (CR) leptons, which can generate bremsstrahlung and inverse Compton (IC) emission on Galactic photon fields, producing a broad spectrum from X-rays to $\gamma$-rays detectable by instruments such as eROSITA and XMM-Newton. For PBHs with masses below $\sim 10^{17}$ g, Hawking evaporation similarly yields low-energy $e^{\pm}$, leading to comparable diffuse emission. Using the first data release from eROSITA and incorporating up-to-date CR propagation and diffusion parameters, we derive new constraints on both scenarios. For sub-GeV DM, we exclude thermally averaged annihilation cross sections in the range $\sim 10^{-27}$$10^{-25}\ \mathrm{cm^{3}/s}$ and decay lifetimes of $\sim 10^{24}$$10^{25}$ s for masses between 1 MeV and 1 GeV, with eROSITA outperforming previous X-ray constraints below $\sim 30$ MeV. For asteroid-mass PBHs, we set new bounds on the DM fraction based on their Hawking-induced emission. Finally, we revisit previous constraints in Ref. [1] and consequently Ref. [2] derived from XMM-Newton, finding that they were approximately four orders of magnitude too stringent due to the use of the instrument’s geometric solid angle rather than its exposure-weighted solid angle. Upon using the exposure-weighted solid angle, we show that the revised XMM-Newton limits are slightly weaker than those from eROSITA.
@@ 1 Introduction
In this work, we address this gap by placing new constraints on light DM and sub-asteroid mass PBHs using eROSITA’s full-sky X-ray maps. We model continuum emission from DM-induced low-energy $e^{\pm}$ and Hawking evaporation from PBHs, deriving upper limits on the annihilation cross section, decay lifetime, and PBH DM fraction across a wide mass range. We then compare these constraints with existing bounds from other probes, demonstrating the complementarity and reach afforded by eROSITA. Additionally, we revisit constraints in Refs. [1, 2] derived from XMM-Newton data [28], showing that applying the geometric solid angle instead of the exposure-weighted solid angle led to flux differences of approximately four orders of magnitude. After addressing this, the XMM-Newton limits are found to be weaker than those from eROSITA, realigning the landscape of X-ray DM constraints. We note that the above discussion above does not apply to Ref. [28] which derives constraints using line-like signals from decaying DM from XMM-Newton.
@@ 3.1 Sub-GeV dark matter annihilation and decay
We restrict our analysis to the channel of direct $e^{+}e^{-}$ production, from a minimum kinematically allowable mass of $m_{e}$ ($2m_{e}$) for annihilation (decay) respectively. In both cases, we take the maximum mass of DM to be $1000$ MeV. We use a 2D axisymmetric Galactic model with a NFW DM profile [36] representing our base case. Other decay channels including the production of pions and muons can also be considered, but the constraints become much weaker than in previous analyses  [1, 2] for the high DM masses needed. The propagation parameters used in this work are obtained from a fit to AMS-02 data [37, 38], and are listed in Table I of Ref. [2], where more of the technical details are given. The interplay between diffusion and energy losses is key at sub-GeV energies. Energy losses in the DRAGON2 code include ionisation, Coulomb, bremsstrahlung, IC, synchrotron and in-flight positron annihilation processes. The spatial and energy resolution employed is $\sim 150$ pc and $5$%.
@@ 4 Constraints from eROSITA data
Finally, in the bottom left panel of Fig. 3, we also report constraints on PBHs in the mass range $10^{14}\ \mathrm{g}\lesssim m_{\mathrm{PBH}}\lesssim 10^{17}\ \mathrm{g}$ from Hawking evaporation of charged particles (orange line). As commented above, these constraints are derived assuming monochromatic and non-rotating PBHs. The orange band again represent uncertainties arising from propagation. Here, we also compare our result from eROSITA with other leading constraints, including the 511 keV line and Voyager 1 (as green lines). In comparison to these probes, eROSITA diffuse background observations do not provide competitive constraints.
@@ 5 Revising XMM-Newton constraints
The DM constraints derived from eROSITA data are significantly weaker than those reported in earlier analyses of XMM-Newton observations [1, 2]. At first glance, this appears counterintuitive. While eROSITA probes lower energy bands than XMM-Newton, where the latter suffers from increased instrumental noise, one would still expect comparable fluxes in the same regions of the sky when interpreted correctly.
We believe this discrepancy arises from a misinterpretation in these previous analyses regarding the solid angle associated with the data presented in Ref. [28]. In particular, earlier works assumed the reported fluxes were distributed over the full geometric solid angle of each ring. However, the data in Ref. [28] are normalised to the exposure-weighted average solid angle $\Omega_{\mathrm{exp}}$, which accounts for the non-uniform exposure across the ring due to the telescope’s observational footprint [^2] A Jupyter notebook tutorial on how to interpret the data is now provided in the original Github repository.. This is defined
@@
Figure 5: Comparison of ring 3 flux for eROSITA (orange points) and XMM-Newton MOS with the use of exposure-weighted solid angle (black points) consistent with observational coverage of the dataset and with direct use of the geometric solid angle (grey points). In order to enable a viable comparison, all fluxes shown here use a 2 mask of the Galactic plane, as provided for the XMM-Newton data.
The discrepancy in solid angle varies between rings, so a full reanalysis is required to derive revised XMM-Newton constraints. Using the same analysis framework as for eROSITA, we compute updated XMM-Newton limits on DM annihilation and decay into $e^{\pm}$, as well as PBH evaporation, shown in Fig. 6. Here we consider energies from 2.5-8 keV, identical to previous analyses. These incorporate propagation uncertainties consistent with those used in the eROSITA analysis. We do not repeat the uncertainty propagation for the DM density distribution or the local DM density as this can be inferred easily from Fig. 4.
@@ 6 Discussion and conclusions
Finally, we reanalysed previous constraints from XMM-Newton data in Ref. [1] and subsequently in Ref. [2], where we identified a discrepancy arising from the treatment of the instrument’s exposure-weighted solid angle. Specifically, an overestimate of the effective sky coverage led to flux values, and therefore constraints, being overstated by approximately four orders of magnitude. After correcting this effect and rederiving the XMM-Newton limits using consistent assumptions, we find that they are weaker than those derived from eROSITA.